The solubility of K2Cr2O7 in water is 125 g/L at 20 °C. A solution is prepared at 20 °C that contains 6.0 grams of K2Cr2O7 in 50. mL of water. This solution is ?

I know the answer is unsatrated but can you explain how and how can I calculate it ?

what I did I took the molarity and * by the volume

which is 6.0*0.05 = 0.3

then I took the mole and* by the formula weight which is

0.3*294=88

so the 88 is less than the solibility which is 125

so it will be undersaturated

does my work correct ?
Answer #1

Apply

the solubility is 125 g/L

then

calcualte concentration of experiment

C = m/V = 6/50 = 0.12 g /mL

change to Liter = 0.12 g / mL * 1000 mL / L = 120 g/L

since 120 < 125, then this is unsaturated